# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Polynomials in arbitrary $n$ of $W^{\mathrm{RW}}_{n,m}(r)$ and $W^{\mathrm{SAW}}_{n,m}(r)$.—Using the enumerated $W^{\mathrm{RW}}_{n,m}(r)$ and $W^{\mathrm{SAW}}_{n,m}(r)$, we derived the polynomials in positive integer $n$ that exactly reproduced the numbers $W^{\mathrm{RW}}_{n,m}(r)$ and $W^{\mathrm{SAW}}_{n,m}(r)$ for $r=n$, $n-2$, $n-4$, $n-6$, $n-8$, and $n-10$. These polynomials enable the calculation of any physical quantity derived from Eq. (2) for $0\leq n-r\leq 10$.
@@
Figure 5: Universal scaling of the internal energy $U_{n}(r,\beta\varepsilon)/\varepsilon$ for (a,b) the DJ model across RW–SAW crossover and (c,d) the ISAW across SAW–NAW crossover. In each panel, points correspond to $n=10,11,\ldots,20$ and $42\leq r\leq n-$ 4; curves represent the four analytic expressions for $U_{n}(r,\beta\varepsilon)/\varepsilon$ obtained from the polynomials of $W^{\mathrm{RW}}_{n,m}(r)$ and $W^{\mathrm{SAW}}_{n,m}(r)$ for $r=n-4$ (gray), $n-6$ (orange), $n-8$ (pink), and $n-10$ (red). In (a,c), $\beta\varepsilon=1$; for a fixed $n$, $U_{n}(r,\beta\varepsilon)/\varepsilon$ monotonically decreases with increasing $r$. In (b,d), $U_{n}(r,\beta\varepsilon)/\varepsilon$, plotted against the scaled variable $(n-r)^{7/4}/n$, collapses onto master curves with a common scaling exponent $7/4$ for a wide range of interaction strengths ($\beta\varepsilon=0$, $1/4$, $1/2$, $1$, $2$, and $4$). The master curves in the log–log plots exhibit slopes close to $1$, indicating an almost linear relation between $U_{n}(r,\beta\varepsilon)/\varepsilon$ and $(n-r)^{7/4}/n$.
Emergence of negative energetic elasticities in lattice polymer chains.—As shown in Fig. 3, the two crossovers interconnecting the RW, SAW, and NAW ($\beta k_{U}=0$) reveal that both the DJ model and ISAW exhibit negative energetic elasticity ($\beta k_{U}<0$) with $|k_{U}/k|$ maximized around $\beta\varepsilon\sim 1$. For $(n,r)=(20,10)$ and the same $\beta\varepsilon$, $k^{\mathrm{DJ}}$ is $4.6$ to $8.7$ times larger than $k^{\mathrm{ISAW}}$ because the configuration space of the RW is larger than that of the SAW. The ratio of $|(k_{U}/k)^{\mathrm{DJ}}|$ to $|(k_{U}/k)^{\mathrm{ISAW}}|$ ranges from $1.1$ to $2.7$, demonstrating the enhanced negative energetic elasticity in the DJ model compared to that in the ISAW.
@@
The DJ model provides an intuitive interpretation of the emergence of negative energetic elasticity. The stiffness at $r=10$ shown in Fig. 3 is calculated from the differences in the free energy at $r=8$, $10$, and $12$. Therefore, to elucidate the underlying mechanism, we consider the $r$-dependence of both entropy and internal energy. Figure 4(a) illustrates two configurations of a coarse-grained polymer chain with soft-core repulsion: a reference state and a stretched state. The energy increases at intersections within the chains [pink triangles in Fig. 4(a)]. The stretched state, which possesses fewer intersections, is thus energetically favorable compared to the reference state. This property holds for the ensemble of chain configurations: states with longer end-to-end distances are energetically favorable; those with shorter end-to-end distances are entropically favorable. This behavior is demonstrated by the monotonic decrease of both the entropy $S_{n}^{\mathrm{DJ}}(r,\beta\varepsilon)$ [Fig.4(b)] and the internal energy $U_{n}^{\mathrm{DJ}}(r,\beta\varepsilon)$ [Fig. 4(c)] with increasing $r$. This interplay between energetic and entropic factors gives rise to negative energetic elasticity.
Scaling in internal energy.—Figures 5(a) (DJ model) and 5(c) (ISAW) show the exact values of $U_{n}(r,\beta\varepsilon)/\varepsilon$ as a function of the chain length $n$ for $n=10,11,\ldots,20$ and $\beta\varepsilon=1$. Here, the end-to-end distance range is $42\leq r\leq n-$ 4. Also shown are four analytic expressions for $U_{n}(r,\beta\varepsilon)/\varepsilon$ corresponding to $r=n-4$, $n-6$, $n-8$, and $n-10$. Remarkably, when plotted against the scaled variable $(n-r)^{7/4}/n$, the data points and analytic expressions for $U_{n}(r,\beta\varepsilon)/\varepsilon$ collapse onto a single master curve for both the DJ model [Fig. 5(b)] and ISAW [Fig. 5(d)] across the entire range of the two crossovers ($\beta\varepsilon=0$, $1/4$, $1/2$, $1$, $2$, and $4$). The master curves exhibit almost linear behavior, suggesting that $U_{n}(r,\beta\varepsilon)/\varepsilon\propto(n-r)^{7/4}/n$ can be used as a practical approximation.
@@
The component $n-r$ appearing in the scaling variable $(n-r)^{7/4}/n$ is the “slack” of the chain, i.e., the number of segments remaining after assigning $r$ segments in the $+x$-direction to reach the endpoint $(r,0,0)$. That a common exponent $7/4$ emerges from the DJ model and the ISAW suggests that this exponent is universal and depends only on the spatial dimensions of the system. Remarkably, the scaling of the internal energy based on the slack $n-r$ is consistent with the scaling of $\tau^{\mathrm{ISAW}}$ found in Ref. [10], which exhibits an exponent of $3/4$. Here, $\tau\equiv[\partial/\partial(\beta\varepsilon)](-k_{U}/k)|_{\beta\varepsilon=0}$ is the first-order coefficient in the Taylor expansion of $-k_{U}/k$ with respect to $\beta\varepsilon$ (see bottom panels of Fig. 3); Ref. [10] uses the notation $T^{\infty}_{U}$ instead of $\tau$. The exponent $3/4$ coincides with the well-established universal critical exponent $\nu=3/4$ of the two-dimensional SAW [40, 41]. This coincidence suggests a possible reduction from three to two dimensions induced by the on-axis constraint on the end-to-end vector (see Sec. S8 in Supplemental Materials of Ref. [10]). Furthermore, the simple relationship between the exponents (i.e., $7/4=3/4+1$) can potentially be explained by introducing a scaling function similar to that of the Widom scaling hypothesis [42, 43]. The analysis of the DJ model under the on-axis constraint presented here is expected to advance the long-standing effort, dating back to the 1970s [14], to connect the critical exponents of the three-dimensional RW and SAW [44].
@@
Conclusion—We elucidated the microscopic origin of negative energetic elasticity in polymer chains using the DJ model and ISAW. Our investigation of the two crossovers among the RW, SAW, and NAW (Fig. 3) revealed that effective soft-repulsive interactions between polymer segments are the origin of the negative energetic elasticity of flexible polymers. The DJ model provides an intuitive interpretation (Fig. 4): longer end-to-end distances are energetically favorable, whereas shorter distances are entropically favorable. We also discovered a universal scaling with an exponent of $7/4$ for the internal energy of both models throughout the crossovers (Fig. 5). This intuitive interpretation of the negative energetic elasticity is expected to serve as a guideline for further theoretical development and material design.

### v2

@@ IV Polynomials in arbitrary $n$ of $W^{\mathrm{RW}}_{n,m}(r)$ and $W^{\mathrm{SAW}}_{n,m}(r)$
Using the enumerated $W^{\mathrm{RW}}_{n,m}(r)$ and $W^{\mathrm{SAW}}_{n,m}(r)$ given in Sec. S1 of the Supplemental Material [27] and Appendix C, we derive the polynomials in positive integer $n$ that exactly reproduce the numbers $W^{\mathrm{RW}}_{n,m}(r)$ and $W^{\mathrm{SAW}}_{n,m}(r)$ for $r=n$, n2n-2, n4n-4, n6n-6, n8n-8, and n10n-10, as shown in Eqs. (13) to (94) and (95) to (106) in Appendixes B and C. These polynomials enable the calculation of any physical quantity derived from Eq. (2) for $0\leq n-r\leq 10$.
To present examples of the polynomials derived using the exact enumeration results, we focus on the polynomials of $W^{\mathrm{RW}}_{n,m}(n-10)$ in $n$ for $m=0$ and $1$, which are given in Eqs. (63) and (64) in Appendix B. To validate these polynomials, we substitute $n=20$ into Eqs. (63) and (64). This substitution yields the values $2625286352$ and $5625406136$, respectively, which are identical to the $m=0$ and $m=1$ values in the $r=10$ column of Table 1. We can extend this validation to $n=21,\dots,26$ by referencing Tables S21 to S23 in the Supplemental Material [27]. By mathematical induction, it is straightforward to prove that Eqs. (63) and (64) hold for all $n\geq 27$.
Figure 3: Emergence of negative energetic elasticity in RW–SAW and SAW–NAW crossovers with $(n,r)=(20,10)$. The DJ model bridges the RW and SAW (left column), and the ISAW bridges the SAW and NAW (right column). The data for the ISAW are obtained from Ref. [10]. The top panels show the dependence of total stiffness $\beta k$ and the entropic contribution to stiffness $\beta k_{S}$ on interaction strength $\beta\varepsilon$. The middle panels show the energetic contribution to stiffness $\beta k_{U}$. The bottom panels show the ratio $k_{U}/k$ and its first-order approximation $\tau\beta\varepsilon$ around $\beta\varepsilon=0$. The maximum of $k_{U}/k$ occurs around $\beta\varepsilon\simeq 1$ for both models.
@@
Figure 5: Universal scaling law of the internal energy $U_{n}(r,\beta\varepsilon)/\varepsilon$ for (a,b) the DJ model across RW–SAW crossover and (c,d) the ISAW across SAW–NAW crossover. In each panel, points correspond to $n=10,11,\ldots,20$ and $2\leq r\leq n-4$; curves represent the four analytic expressions for $U_{n}(r,\beta\varepsilon)/\varepsilon$ obtained from the polynomials of $W^{\mathrm{RW}}_{n,m}(r)$ and $W^{\mathrm{SAW}}_{n,m}(r)$ (see Appendixes B and C) for r=n4r=n-4 (gray), n6n-6 (orange), n8n-8 (pink), and n10n-10 (red). In (a,c), $\beta\varepsilon=1$; for a fixed $n$, $U_{n}(r,\beta\varepsilon)/\varepsilon$ monotonically decreases with increasing $r$. In (b,d), $U_{n}(r,\beta\varepsilon)/\varepsilon$, plotted against the scaled variable $(n-r)^{7/4}/n$, collapses onto master curves with a common scaling exponent $7/4$ for a wide range of interaction strengths ($\beta\varepsilon=0$, $1/4$, $1/2$, 11, 22, and $4$). The master curves show a approximately linear behavior in the small $(n-r)^{7/4}/n$ regime.
V Emergence of negative energetic elasticities in lattice polymer chains
Figure 3 shows $\beta k$, $\beta k_{S}$, $\beta k_{U}$, $k_{U}/k$, and $\tau\beta\varepsilon$ for the DJ model and ISAW for $(n,r)=(20,10)$. The left column of Fig. 3 presents these values for the DJ model, calculated using Table 1, which enables the computation of the differences $\varDelta W_{n,m}(r)$ and $\varDelta^{2}W_{n,m}(r)$ at $r=10$ in Eq. (6). The two crossovers interconnecting the RW, SAW, and NAW ($\beta k_{U}=0$) reveal that both the DJ model and ISAW exhibit negative energetic elasticity ($\beta k_{U}<0$) with $|k_{U}/k|$ maximized around $\beta\varepsilon\sim 1$. For $(n,r)=(20,10)$ and the same $\beta\varepsilon$, $k^{\mathrm{DJ}}$ is $4.6$ to $8.7$ times larger than $k^{\mathrm{ISAW}}$ because the configuration space of the RW is larger than that of the SAW. The ratio of $|(k_{U}/k)^{\mathrm{DJ}}|$ to $|(k_{U}/k)^{\mathrm{ISAW}}|$ ranges from $1.1$ to $2.7$, demonstrating the enhanced negative energetic elasticity in the DJ model compared to that in the ISAW.
@@ V Emergence of negative energetic elasticities in lattice polymer chains
The DJ model suggests an intuitive interpretation of the emergence of negative energetic elasticity in polymer chains. The stiffness at $r=10$ shown in Fig. 3 is calculated from the differences in the free energy at $r=8$, $10$, and $12$. Therefore, to elucidate the underlying mechanism, we consider the $r$-dependence of both entropy and internal energy. Figure 4(a) illustrates two configurations of a coarse-grained polymer chain with soft-core repulsion: a reference state and a stretched state. The energy increases at intersections within the chains [pink triangles in Fig. 4(a)]. The stretched state, which possesses fewer intersections, is thus energetically favorable compared to the reference state. This property holds for the ensemble of chain configurations: states with longer end-to-end distances are energetically favorable; those with shorter end-to-end distances are entropically favorable. This behavior is demonstrated by the monotonic decrease of both the entropy $S_{n}^{\mathrm{DJ}}(r,\beta\varepsilon)$ [Fig.4(b)] and the internal energy $U_{n}^{\mathrm{DJ}}(r,\beta\varepsilon)$ [Fig. 4(c)] with increasing $r$. This interplay between energetic and entropic factors gives rise to negative energetic elasticity.
The insights from the DJ model can be extended to understand the origin of negative energetic elasticity in gel networks, attributed to the same underlying mechanism. In Fig. 4(a), the polymer chain can be considered a subchain between crosslinks in polymer networks, with $r_{\mathrm{ref}}$ corresponding to the average distance between crosslinks at the as-prepared state. Notably, in gels synthesized by end-linking star polymers, where negative energetic elasticity has been observed experimentally [4, 5, 6, 7], molecular dynamics simulations [41] and gel fracture experiments [42] (see the section “Failure of Kuhn’s model for fracture” therein) demonstrate that $r_{\mathrm{ref}}$ is a decreasing function of polymer mass concentration. At lower concentrations, a larger $r_{\mathrm{ref}}$ is expected, and the gel becomes unstable below a certain concentration, as confirmed by observations of gel–gel phase separation [43]. Since $r_{\mathrm{ref}}$ depends on concentration, the monotonic decrease of both $S_{n}^{\mathrm{DJ}}(r,\beta\varepsilon)$ and $U_{n}^{\mathrm{DJ}}(r,\beta\varepsilon)$ with increasing $r$ [Figs. 4(b) and 4(c)] becomes crucial for understanding the origin of negative energetic elasticity in gel networks.
VI Scaling law in internal energy
Figures 5(a) (DJ model) and 5(c) (ISAW) show the exact values of $U_{n}(r,\beta\varepsilon)/\varepsilon$ as a function of the chain length $n$ for $n=10,11,\ldots,20$ at $\beta\varepsilon=1$, with the end-to-end distance $2\leq r\leq n-4$. Also shown are four analytic expressions for $U_{n}(r,\beta\varepsilon)/\varepsilon$ corresponding to r=n4r=n-4, n6n-6, n8n-8, and n10n-10. These analytic expressions for $U_{n}(r,\beta\varepsilon)/\varepsilon$ are derived by using Eqs. (2) and (4), inserting the polynomials for $W^{\mathrm{RW}}_{n,m}(r)$ presented in Sec. IV. As an example, we present the analytic expressions of $U^{\mathrm{DJ}}_{n}(n-4,\beta\varepsilon)/\varepsilon$ (valid for $n\geq 5$), which corresponds to the gray curves in Figs. 5(a) and 5(b):
$U^{\mathrm{DJ}}_{n}(n-4,\beta\varepsilon)/\varepsilon=\frac{\displaystyle\sum_{m=1}^{6}m\,W^{\mathrm{RW}}_{n,m}(n-4)\,e^{-\beta\varepsilon m}}{\displaystyle\sum_{m=0}^{6}W^{\mathrm{RW}}_{n,m}(n-4)\,e^{-\beta\varepsilon m}},$ (9)
with the polynomials $W^{\mathrm{RW}}_{n,m}(n-4)$ given by
$W^{\mathrm{RW}}_{n,0}(n-4)$ $=$ $\frac{1}{2}(3n^{4}-34n^{3}+153n^{2}-322n+248),$
$W^{\mathrm{RW}}_{n,1}(n-4)$ $=$ $4(2n^{3}-15n^{2}+31n-6),$
$W^{\mathrm{RW}}_{n,2}(n-4)$ $=$ $2n^{3}-10n^{2}+36n-81,$
$W^{\mathrm{RW}}_{n,3}(n-4)$ $=$ $4n^{2}-22n+50,$
$W^{\mathrm{RW}}_{n,4}(n-4)$ $=$ $\frac{1}{2}(n^{2}+29n-110),$
$W^{\mathrm{RW}}_{n,5}(n-4)$ $=$ $2(n-5),$
$W^{\mathrm{RW}}_{n,6}(n-4)$ $=$ $n-4,$
$W^{\mathrm{RW}}_{n,m}(n-4)$ $=$ $0\qquad(\textrm{for}\;\;m\geq 7).$
These polynomials are taken from Eqs. (19) to (26) in Appendix B.
A remarkable scaling law emerges when the data points and analytic expressions for $U_{n}(r,\beta\varepsilon)/\varepsilon$ are plotted against the scaled variable $(n-r)^{7/4}/n$. For both the DJ model [Fig. 5(b)] and ISAW [Fig. 5(d)], all data points and analytic expressions for $U_{n}(r,\beta\varepsilon)/\varepsilon$ collapse onto a single master curve across the entire range of $\beta\varepsilon$ (0, $1/4$, $1/2$, 11, 22, and $4$). This scaling law persists even for off-axis end-to-end vectors along $(1,1,0)$ and $(1,1,1)$, as shown in Appendix D. The component nrn-r appearing in the scaling variable $(n-r)^{7/4}/n$ is the “slack” of the chain, i.e., the number of segments remaining after assigning $r$ segments in the positive $x$ direction to reach the endpoint $(r,0,0)$. Maclaurin expansions of the analytic expressions for $U_{n}(n-r,\beta\varepsilon)/\varepsilon$ at nr=4,6,8n-r=4,6,8, and $10$ with respect to $1/n$ yield linear leading-order terms, explaining the observed linear behavior at small $(n-r)^{7/4}/n$.
The emergence of a common exponent $7/4$ in both the DJ model and the ISAW suggests its universality governed by spatial dimensionality, independent of microscopic model details. Remarkably, this scaling law for the internal energy based on the slack nrn-r is consistent with the scaling behavior of $\tau^{\mathrm{ISAW}}$ reported in Ref. [10], which exhibits an exponent of $3/4$. Here,
$\tau\equiv-\frac{\partial}{\partial(\beta\varepsilon)}\left.\left(\frac{k_{U}}{k}\right)\right|_{\beta\varepsilon=0}$ (10)
is the first-order coefficient in the Maclaurin series of $k_{U}/k$ with respect to $\beta\varepsilon$ (see bottom panels of Fig. 3). Reference [10] denotes this coefficient as $T^{\infty}_{U}$ instead of $\tau$.
@@ VI Scaling law in internal energy
The exponent $3/4$ coincides with the well-established universal critical exponent $\nu=3/4$ of the two-dimensional SAW [44, 45]. This coincidence suggests a possible reduction from three to two dimensions induced by the on-axis constraint on the end-to-end vector (see Sec. S8 in the Supplemental Material of Ref. [10]). Furthermore, the simple relationship between the exponents (i.e., $7/4=3/4+1$) can potentially be explained by introducing a scaling function similar to that of the Widom scaling hypothesis [46, 47]. The analysis of the DJ model presented here is expected to advance the long-standing effort, dating back to the 1970s [14], to connect the critical exponents of the three-dimensional RW and SAW [48].
We note that this common scaling exponent does not imply that the RW and SAW belong to the same universality class. RW (Gaussian) and SAW universality classes remain distinct with different critical exponents ($\nu=0.5$ for RW and $\nu\simeq 0.588$ for SAW in three dimensions) that govern the scaling of spatial extent with chain length. The $7/4$ exponent identified here relates specifically to the internal energy under stretching, not to the critical exponents that define these universality classes.
VII Conclusion
@@
We elucidated the microscopic origin of negative energetic elasticity in polymer chains using the DJ model and ISAW. Our investigation of the two crossovers among the RW, SAW, and NAW (Fig. 3) revealed that effective soft-repulsive interactions between polymer segments are the origin of the negative energetic elasticity of flexible polymers. The DJ model provides an intuitive interpretation (Fig. 4): longer end-to-end distances are energetically favorable, whereas shorter distances are entropically favorable. We also discovered a universal scaling law with an exponent of $7/4$ for the internal energy of both models throughout the crossovers (Fig. 5). This intuitive interpretation of the negative energetic elasticity, together with the universal scaling law, is expected to serve as a guideline for further theoretical development and material design.
These findings have significant implications for understanding the elasticity of polymer gel networks. Our analysis suggests that negative energetic elasticity is a universal property of polymer systems with effective soft-repulsive segment interactions, irrespective of their composition or architecture. This understanding provides a fundamental framework for explaining the recently discovered negative energetic elasticity in various gels and opens new avenues for designing gel materials with tailored elastic properties.